# second generation 3G third generation

### v1

@@ III-A PARAFAC-Based Receiver
where $\overline{\boldsymbol{G}}=\Big[\boldsymbol{g}_{1}\;\dots\;\boldsymbol{g}_{I}\Big]^{\text{T}}\in\mathbb{C}^{I\times M_{T}N}$. Note that $\boldsymbol{\mathcal{Z}}$ matches a third-order PARAFAC tensor model [Harshman]. In Figure 2, we illustrate the procedures run until this point. It can be noted that the received signal model $\boldsymbol{Y}_{i,k}$ is a PARATUCK(2,4), which is a fourth-order tensor $\boldsymbol{\mathcal{Y}}$. Concatenating this tensor via the vectorized version related to the $3$-mode $K$, for each frame we have a third-order tensor $\boldsymbol{\mathcal{Y}}_{i}\in\mathbb{C}^{M_{R}\times T\times K}$ that matches a third-order TUCKER model, $i=1,\ldots,I$. However, the $3$-mode unfolding transposed of $\boldsymbol{\mathcal{Y}}_{i}$ can be seen as the $i$-th frontal slice of a tensor $\boldsymbol{\mathcal{Z}}\in\mathbb{C}^{M_{R}T\times K\times I}$ that fits a PARAFAC model where the factor matrices are structured. Let us recall that $\boldsymbol{\Omega}$ is the Kronecker product between the data symbols $\boldsymbol{X}$ and the equivalent channel $\boldsymbol{HS}$ while $\boldsymbol{\Psi}$ is the Khatri-Rao (column-wise Kronecker) product between the rotation matrix $\boldsymbol{P}$ and the coding matrix $\boldsymbol{W}$. In this sense, the following correspondence can be established:
Figure 3: Third-order PARAFAC tensor model: The received signal at $(14)$ fits a fourth-order PARATUCK tensor model. From the third mode unfolding of the tensor $\boldsymbol{\mathcal{Y}}_{i,k}$, we derive a third-order PARAFAC model, with $\boldsymbol{\Omega}$, $\boldsymbol{\Psi}$ and $\overline{\boldsymbol{G}}$ as factor matrices.
$\Big(\boldsymbol{A},\;\boldsymbol{B},\;\boldsymbol{C}\Big)$ $\leftrightarrow\Big(\boldsymbol{\Omega},\;\boldsymbol{\Psi},\;\overline{\boldsymbol{G}}\Big)$ (23)
@@
To estimate the factor matrix $\boldsymbol{\Omega}$ and the channel $\overline{\boldsymbol{G}}$, we must solve the following optimization problem
@@
Note that (24) is a nonlinear problem due to the presence of two unknown matrices $(\overline{\boldsymbol{G}}\;\text{and}\;\boldsymbol{\Omega})$. To address this issue, we can derive sub-optimization linear programs that hold certain matrices constant. For instance, by fixing the channel matrix $\overline{\boldsymbol{G}}$, the problem becomes linear in the factor matrix $\boldsymbol{\Omega}$. Similarly, fixing $\boldsymbol{\Omega}$ we have a linear problem in $\overline{\boldsymbol{G}}$. Consequently, the optimization problem in (24) can be decomposed into sub-optimization problems, which can then be solved iteratively. This is the underlying premise of the alternating least squares (ALS) algorithm. On the other hand, since we are interested in solving only for $\boldsymbol{\Omega}$ and $\overline{\boldsymbol{G}}$, the algorithm is defined as bilinear alternating least squares (BALS). For this, we shall make use of $\boldsymbol{\mathcal{Z}}]_{(1)}$ and $\boldsymbol{\mathcal{Z}}]_{(3)}$, which from (20), are defined as
(1) $=\Big[[\boldsymbol{\mathcal{Y}}_{1}]_{(3)}^{\text{T}}\;[\boldsymbol{\mathcal{Y}}_{2}]_{(3)}^{\text{T}}\;\dots\;[\boldsymbol{\mathcal{Y}}_{I}]_{(3)}^{\text{T}}\Big]$ (25)
$=\boldsymbol{\Omega}\Big[\text{diag}(\boldsymbol{g}_{1})\boldsymbol{\Psi}^{\text{T}}\;\text{diag}(\boldsymbol{g}_{2})\boldsymbol{\Psi}^{\text{T}}\;\dots\;\text{diag}(\boldsymbol{g}_{I})\boldsymbol{\Psi}^{\text{T}}\Big]$
$=\boldsymbol{\Omega}\Big(\overline{\boldsymbol{G}}\diamond\boldsymbol{\Psi}\Big)^{\text{T}}\;\in\mathbb{C}^{TM_{R}\times IK}\;,$
@@
(3) $=\Big[\text{vec}([\boldsymbol{\mathcal{Y}}_{1}]_{(3)}^{\text{T}})\;\text{vec}([\boldsymbol{\mathcal{Y}}_{2}]_{(3)}^{\text{T}})\;\dots\;\text{vec}([\boldsymbol{\mathcal{Y}}_{I}]_{(3)}^{\text{T}})\Big]^{\text{T}}$ (26)
$=\Big[\big(\boldsymbol{\Psi}\diamond\boldsymbol{\Omega}\big)[\boldsymbol{g}_{1}\;\boldsymbol{g}_{2}\;\dots\;\boldsymbol{g}_{I}]\Big]^{\text{T}}$
$=\overline{\boldsymbol{G}}\big(\boldsymbol{\Psi}\diamond\boldsymbol{\Omega}\big)^{\text{T}}\in\mathbb{C}^{I\times KTM_{R}}.$
By exploiting the first and third mode unfoldings, as described in equations (25) and (26), respectively, we propose an iterative solution to estimate the structured matrix $\boldsymbol{\Omega}$ and the channel matrix $\overline{\boldsymbol{G}}$. It is important to note that after estimating $\boldsymbol{\Omega}$, a Kronecker factorization problem must be solved to recover the data symbol matrix $\boldsymbol{X}$ and the BS-IRS channel $\boldsymbol{H}$. This results in a two-stage solution: first, we solve the PARAFAC problems using the BALS algorithm, followed by a Kronecker factorization problem. Due to this combined approach, we have titled this algorithm PAKRON (Algorithm 1).
@@
input : Received signal $\boldsymbol{\mathcal{Z}}$
output : $\hat{\boldsymbol{X}}$, $\hat{\boldsymbol{H}}$ and $\overline{\boldsymbol{G}}$
@@ III-A1 PAKRON receiver (Stage I)
By exploiting the matrix unfoldings (25) and (26), we derive the BALS algorithm. This algorithm estimates the matrices $\boldsymbol{\Omega}$ and $\overline{\boldsymbol{G}}$ in an alternating manner by iteratively optimizing the following two cost functions:
$\hat{\boldsymbol{\Omega}}=\underset{\overline{\boldsymbol{G}}}{\arg\min}\,\,\left\|[\boldsymbol{\mathcal{Y}}]_{(1)}-\boldsymbol{\Omega}\Big(\overline{\boldsymbol{G}}\diamond\boldsymbol{\Psi}\Big)^{\textrm{T}}\right\|_{F}^{2},$ (27)
$\hat{\overline{\boldsymbol{G}}}^{\text{T}}=\underset{\boldsymbol{\Omega}}{\arg\min}\,\,\left\|[\boldsymbol{\mathcal{Y}}]_{(3)}-\overline{\boldsymbol{G}}^{\text{T}}\Big(\boldsymbol{\Psi}\diamond\boldsymbol{\Omega}\Big)^{\textrm{T}}\right\|_{F}^{2},$ (28)
@@
$\hat{\boldsymbol{\Omega}}=[\boldsymbol{\mathcal{Y}}]_{(1)}\Big[\Big(\overline{\boldsymbol{G}}\diamond\boldsymbol{\Psi}\Big)^{\textrm{T}}\Big]^{\dagger},$ (29)
$\hat{\overline{\boldsymbol{G}}}=[\boldsymbol{\mathcal{Y}}]_{(3)}\Big[\Big(\boldsymbol{\Psi}\diamond\boldsymbol{\Omega}\Big)^{\textrm{T}}\Big]^{\dagger}.$ (30)
@@
input : Received signal $\boldsymbol{\mathcal{Z}}$
output : $\hat{\boldsymbol{\Omega}}$, $\hat{\overline{\boldsymbol{G}}}$
1exto $i=0$ initialize $\hat{\overline{\boldsymbol{G}}}_{(i=0)}$ begin
@@
    Find a least squares estimate of $\hat{\boldsymbol{\Omega}}i$:
  • $\hat{\boldsymbol{\Omega}}_{i}=\big[\boldsymbol{\mathcal{Y}}\big]_{(1)}\big[\big(\overline{\boldsymbol{G}}_{(i-1)}\diamond\boldsymbol{\Psi}\big)^{\textrm{T}}\big]^{\dagger}$
@@
    Find a least squares estimate of $\hat{\overline{\boldsymbol{G}}}_{i}$:
    • $\hat{\overline{\boldsymbol{G}}}_{i}=\big[\boldsymbol{\mathcal{Y}}\big]_{(3)}\big[\big(\boldsymbol{\Psi}\diamond\hat{\boldsymbol{\Omega}}_{i}\big)^{\textrm{T}}\big]^{\dagger}$
@@
    Repeat steps $1$ to $2$ until convergence.
    end
@@
    The convergence of Algorithm 4 is declared when $\|e_{(i)}-e_{(i-1)}\|\leq\delta$, where $e_{(i)}=\|\mathcal{Y}-\hat{\mathcal{Y}}_{(i)}\|_{F}^{2}$ denotes the reconstruction error computed at the $i$-th iteration. $\delta$ is a predefined threshold parameter, and $\hat{\mathcal{Y}}_{(i)}=[\hat{\overline{\boldsymbol{G}}}_{(i)},\hat{\boldsymbol{\Omega}}_{(i)},\boldsymbol{\Psi}]$ is the reconstructed PARAFAC model at the $i$-th iteration. From Stage I, we obtain an estimate of $\boldsymbol{\Omega}$, as defined in equation (18), which is subsequently used in Stage II.
@@ III-A2 PAKRON receiver (Stage II)
    In Stage II, the estimation of the data symbol matrix $\boldsymbol{X}$ and the RIS-BS channel $\boldsymbol{H}$ is achieved by solving the following optimization problem:
    $(\hat{\boldsymbol{X}},\hat{\boldsymbol{H}})=\underset{\boldsymbol{X},\boldsymbol{H}}{\arg\min}\|\boldsymbol{\Delta}-\boldsymbol{X}\otimes\boldsymbol{H}\|_{\text{F}}^{2}.$ (31)
    where $\boldsymbol{\Delta}=\hat{\boldsymbol{\Omega}}(\boldsymbol{I}_{M_{T}}\otimes\boldsymbol{S}^{\text{H}}).$ The matrix $\overline{\boldsymbol{Z}}$ can be approximated as a rank one matrix $\boldsymbol{\Delta}$, such that:
    $\begin{split}\tilde{\boldsymbol{\Delta}}&=\left[\begin{array}[]{ccc}x_{1,1}\textrm{vec}(\boldsymbol{H})&\dots&x_{T,UL}\textrm{vec}(\boldsymbol{H})\end{array}\right]^{\textrm{T}}\\ &=\textrm{vec}(\boldsymbol{X})\textrm{vec}(\boldsymbol{H})^{\textrm{T}}\in\mathbb{C}^{TM_{T}\times M_{R}N}.\end{split}$ (32)
@@
    $(\hat{\boldsymbol{X}},\hat{\boldsymbol{H}})=\underset{\boldsymbol{X},\boldsymbol{H}}{\min}\|\widetilde{\boldsymbol{Z}}-\boldsymbol{x}\boldsymbol{h}^{\textrm{T}}\|_{\text{F}},$ (33)
    where $\boldsymbol{x}=\textrm{vec}(\boldsymbol{X})$ and $\boldsymbol{h}=\textrm{vec}(\boldsymbol{H})$. According to [Eckart_36], the optimal solution for estimating the vectorized version of the RIS-BS channel and the data symbol matrix is obtained by taking the left and right singular vectors of $\tilde{\boldsymbol{\Delta}}$, as illustrated in Algorithm 3.
@@
    input : $\hat{\boldsymbol{\Omega}}$
    output : $\hat{\boldsymbol{X}}$ and $\hat{\boldsymbol{H}}$
@@
      Construct the rank-one matrix $\tilde{\boldsymbol{\Delta}}\in\mathbb{C}^{TM_{T}\times M_{R}N}$
@@
    $\hat{\boldsymbol{x}}\longleftarrow\sqrt{\sigma_{1}}\boldsymbol{u}_{1}$; $\hat{\boldsymbol{h}}\longleftarrow\sqrt{\sigma_{1}}\boldsymbol{v}_{1}^{\ast}$
@@
    Reconstruct $\hat{\boldsymbol{X}}$ and $\hat{\boldsymbol{H}}$ by unvec $\hat{\boldsymbol{x}}$ and $\hat{\boldsymbol{h}}$ 4.   Remove the scaling ambiguities of $\hat{\boldsymbol{X}}$ and $\hat{\boldsymbol{H}}$. end

### v2

@@ III-A PARAFAC-Based Receiver
where $\overline{\bm{G}}=\Big[\bm{g}_{1}\;\dots\;\bm{g}_{I}\Big]^{\text{T}}\in\mathbb{C}^{I\times M_{T}N}$. Note that $\bm{\mathcal{Z}}$ matches a third-order PARAFAC tensor model [Harshman]. In Figure 2, we illustrate the procedures run until this point. It can be noted that the received signal model $\bm{Y}_{i,k}$ is a PARATUCK(2,4), which is a fourth-order tensor $\bm{\mathcal{Y}}$. Concatenating this tensor via the vectorized version related to the $3$-mode $K$, for each frame we have a third-order tensor $\bm{\mathcal{Y}}_{i}\in\mathbb{C}^{M_{R}\times T\times K}$ that matches a third-order TUCKER model, $i=1,\ldots,I$. However, the $3$-mode unfolding transposed of $\bm{\mathcal{Y}}_{i}$ can be seen as the $i$-th frontal slice of a tensor $\bm{\mathcal{Z}}\in\mathbb{C}^{M_{R}T\times K\times I}$ that fits a PARAFAC model where the factor matrices are structured. Let us recall that $\bm{\Omega}$ is the Kronecker product between the data symbols $\bm{X}$ and the equivalent channel $\bm{HS}$ while $\bm{\Psi}$ is the Khatri-Rao (column-wise Kronecker) product between the rotation matrix $\bm{P}$ and the coding matrix $\bm{W}$. In this sense, the following correspondence can be established:
Figure 3: Third-order PARAFAC tensor model: The received signal at $(14)$ fits a fourth-order PARATUCK tensor model. From the third mode unfolding of the tensor $\bm{\mathcal{Y}}_{i,k}$, we derive a third-order PARAFAC model, with $\bm{\Omega}$, $\bm{\Psi}$ and $\overline{\bm{G}}$ as factor matrices.
$\Big(\bm{A},\;\bm{B},\;\bm{C}\Big)$ $\leftrightarrow\Big(\bm{\Omega},\;\bm{\Psi},\;\overline{\bm{G}}\Big)$ (23)
@@
To estimate the factor matrix $\bm{\Omega}$ and the channel $\overline{\bm{G}}$, we must solve the following optimization problem
@@
Note that (24) is a nonlinear problem due to the presence of two unknown matrices $(\overline{\bm{G}}\;\text{and}\;\bm{\Omega})$. To address this issue, we can derive sub-optimization linear programs that hold certain matrices constant. For instance, by fixing the channel matrix $\overline{\bm{G}}$, the problem becomes linear in the factor matrix $\bm{\Omega}$. Similarly, fixing $\bm{\Omega}$ we have a linear problem in $\overline{\bm{G}}$. Consequently, the optimization problem in (24) can be decomposed into sub-optimization problems, which can then be solved iteratively. This is the underlying premise of the alternating least squares (ALS) algorithm. On the other hand, since we are interested in solving only for $\bm{\Omega}$ and $\overline{\bm{G}}$, the algorithm is defined as bilinear alternating least squares (BALS). For this, we shall make use of $\bm{\mathcal{Z}}]_{(1)}$ and $\bm{\mathcal{Z}}]_{(3)}$, which from (20), are defined as
(1) $=\Big[[\bm{\mathcal{Y}}_{1}]_{(3)}^{\text{T}}\;[\bm{\mathcal{Y}}_{2}]_{(3)}^{\text{T}}\;\dots\;[\bm{\mathcal{Y}}_{I}]_{(3)}^{\text{T}}\Big]$ (25)
$=\bm{\Omega}\Big[\text{diag}(\bm{g}_{1})\bm{\Psi}^{\text{T}}\;\text{diag}(\bm{g}_{2})\bm{\Psi}^{\text{T}}\;\dots\;\text{diag}(\bm{g}_{I})\bm{\Psi}^{\text{T}}\Big]$
$=\bm{\Omega}\Big(\overline{\bm{G}}\diamond\bm{\Psi}\Big)^{\text{T}}\;\in\mathbb{C}^{TM_{R}\times IK}\;,$
@@
(3) $=\Big[\text{vec}([\bm{\mathcal{Y}}_{1}]_{(3)}^{\text{T}})\;\text{vec}([\bm{\mathcal{Y}}_{2}]_{(3)}^{\text{T}})\;\dots\;\text{vec}([\bm{\mathcal{Y}}_{I}]_{(3)}^{\text{T}})\Big]^{\text{T}}$ (26)
$=\Big[\big(\bm{\Psi}\diamond\bm{\Omega}\big)[\bm{g}_{1}\;\bm{g}_{2}\;\dots\;\bm{g}_{I}]\Big]^{\text{T}}$
$=\overline{\bm{G}}\big(\bm{\Psi}\diamond\bm{\Omega}\big)^{\text{T}}\in\mathbb{C}^{I\times KTM_{R}}.$
By exploiting the first and third mode unfoldings, as described in equations (25) and (26), respectively, we propose an iterative solution to estimate the structured matrix $\bm{\Omega}$ and the channel matrix $\overline{\bm{G}}$. It is important to note that after estimating $\bm{\Omega}$, a Kronecker factorization problem must be solved to recover the data symbol matrix $\bm{X}$ and the BS-IRS channel $\bm{H}$. This results in a two-stage solution: first, we solve the PARAFAC problems using the BALS algorithm, followed by a Kronecker factorization problem. Due to this combined approach, we have titled this algorithm PAKRON (Algorithm 1).
@@
input : Received signal $\bm{\mathcal{Z}}$
output : $\hat{\bm{X}}$, $\hat{\bm{H}}$ and $\overline{\bm{G}}$
@@ III-A1 PAKRON receiver (Stage I)
By exploiting the matrix unfoldings (25) and (26), we derive the BALS algorithm. This algorithm estimates the matrices $\bm{\Omega}$ and $\overline{\bm{G}}$ in an alternating manner by iteratively optimizing the following two cost functions:
$\hat{\bm{\Omega}}=\underset{\overline{\bm{G}}}{\arg\min}\,\,\left\|[\bm{\mathcal{Y}}]_{(1)}-\bm{\Omega}\Big(\overline{\bm{G}}\diamond\bm{\Psi}\Big)^{\textrm{T}}\right\|_{F}^{2},$ (27)
$\hat{\overline{\bm{G}}}^{\text{T}}=\underset{\bm{\Omega}}{\arg\min}\,\,\left\|[\bm{\mathcal{Y}}]_{(3)}-\overline{\bm{G}}^{\text{T}}\Big(\bm{\Psi}\diamond\bm{\Omega}\Big)^{\textrm{T}}\right\|_{F}^{2},$ (28)
@@
$\hat{\bm{\Omega}}=[\bm{\mathcal{Y}}]_{(1)}\Big[\Big(\overline{\bm{G}}\diamond\bm{\Psi}\Big)^{\textrm{T}}\Big]^{\dagger},$ (29)
$\hat{\overline{\bm{G}}}=[\bm{\mathcal{Y}}]_{(3)}\Big[\Big(\bm{\Psi}\diamond\bm{\Omega}\Big)^{\textrm{T}}\Big]^{\dagger}.$ (30)
@@
input : Received signal $\bm{\mathcal{Z}}$
output : $\hat{\bm{\Omega}}$, $\hat{\overline{\bm{G}}}$
1exto $i=0$ initialize $\hat{\overline{\bm{G}}}_{(i=0)}$ begin
@@
    Find a least squares estimate of $\hat{\bm{\Omega}}i$:
  • $\hat{\bm{\Omega}}_{i}=\big[\bm{\mathcal{Y}}\big]_{(1)}\big[\big(\overline{\bm{G}}_{(i-1)}\diamond\bm{\Psi}\big)^{\textrm{T}}\big]^{\dagger}$
@@
    Find a least squares estimate of $\hat{\overline{\bm{G}}}_{i}$:
    • $\hat{\overline{\bm{G}}}_{i}=\big[\bm{\mathcal{Y}}\big]_{(3)}\big[\big(\bm{\Psi}\diamond\hat{\bm{\Omega}}_{i}\big)^{\textrm{T}}\big]^{\dagger}$
@@
    Repeat steps $1$ to $2$ until convergence. end
@@
    The convergence of Algorithm 4 is declared when $\|e_{(i)}-e_{(i-1)}\|\leq\delta$, where $e_{(i)}=\|\mathcal{Y}-\hat{\mathcal{Y}}_{(i)}\|_{F}^{2}$ denotes the reconstruction error computed at the $i$-th iteration. $\delta$ is a predefined threshold parameter, and $\hat{\mathcal{Y}}_{(i)}=[\hat{\overline{\bm{G}}}_{(i)},\hat{\bm{\Omega}}_{(i)},\bm{\Psi}]$ is the reconstructed PARAFAC model at the $i$-th iteration. From Stage I, we obtain an estimate of $\bm{\Omega}$, as defined in equation (18), which is subsequently used in Stage II.
@@ III-A2 PAKRON receiver (Stage II)
    In Stage II, the estimation of the data symbol matrix $\bm{X}$ and the RIS-BS channel $\bm{H}$ is achieved by solving the following optimization problem:
    $(\hat{\bm{X}},\hat{\bm{H}})=\underset{\bm{X},\bm{H}}{\arg\min}\|\bm{\Delta}-\bm{X}\otimes\bm{H}\|_{\text{F}}^{2}.$ (31)
    where $\bm{\Delta}=\hat{\bm{\Omega}}(\bm{I}_{M_{T}}\otimes\bm{S}^{\text{H}}).$ The matrix $\overline{\bm{Z}}$ can be approximated as a rank one matrix $\bm{\Delta}$, such that:
    $\begin{split}\tilde{\bm{\Delta}}&=\left[\begin{array}[]{ccc}x_{1,1}\textrm{vec}(\bm{H})&\dots&x_{T,UL}\textrm{vec}(\bm{H})\end{array}\right]^{\textrm{T}}\\ &=\textrm{vec}(\bm{X})\textrm{vec}(\bm{H})^{\textrm{T}}\in\mathbb{C}^{TM_{T}\times M_{R}N}.\end{split}$ (32)
@@
    $(\hat{\bm{X}},\hat{\bm{H}})=\underset{\bm{X},\bm{H}}{\min}\|\widetilde{\bm{Z}}-\bm{x}\bm{h}^{\textrm{T}}\|_{\text{F}},$ (33)
    where $\bm{x}=\textrm{vec}(\bm{X})$ and $\bm{h}=\textrm{vec}(\bm{H})$. According to [Eckart_36], the optimal solution for estimating the vectorized version of the RIS-BS channel and the data symbol matrix is obtained by taking the left and right singular vectors of $\tilde{\bm{\Delta}}$, as illustrated in Algorithm 3.
@@
    input : $\hat{\bm{\Omega}}$
    output : $\hat{\bm{X}}$ and $\hat{\bm{H}}$
@@
      Construct the rank-one matrix $\tilde{\bm{\Delta}}\in\mathbb{C}^{TM_{T}\times M_{R}N}$
@@
    $\hat{\bm{x}}\longleftarrow\sqrt{\sigma_{1}}\bm{u}_{1}$; $\hat{\bm{h}}\longleftarrow\sqrt{\sigma_{1}}\bm{v}_{1}^{\ast}$
@@
    Reconstruct $\hat{\bm{X}}$ and $\hat{\bm{H}}$ by unvec $\hat{\bm{x}}$ and $\hat{\bm{h}}$
  • 4.
    Remove the scaling ambiguities of $\hat{\bm{X}}$ and $\hat{\bm{H}}$. end